OpenStudy (anonymous):

n and a are integers. Explain why (n^2 - a^2) - (n - a)^2 is always an even integer.

OpenStudy (anonymous):

u can write (n^2 - a^2) as (n-a)(n+a) (n-a)(n+a)-(n - a)^2 (n-a)[n+a-n+a] (n-a)2a 2[a(n-a)] = multiple of 2 always soeven

OpenStudy (anonymous):

used formula a2-b2=(a-b)(a+b)